OpenStudy (anonymous):

I am trying to figure out how to formulate problems like such: You travel 13 miles north, 10 miles east and 4 miles south; how far are you from where you started?

OpenStudy (dumbcow):

think of coord points start at (0,0) 13 north (0,13) 10 east (10,13) 4 south (10,9) distance = sqrt(10^2 + 9^2) = sqrt(181) = 13.45